# Energetic particles accelerated via turbulent magnetic reconnection in protoplanetary discs – I: ionisation rates

### v1

@@ 3.3 Normalisation of the energy distribution from the available energy of the turbulent accretion
We normalise the non-thermal distribution by assuming that the power sustaining it, $\dot{U}_{\rm nt}$, is a fixed fraction $\kappa$ of the local viscous dissipation per unit volume, $D/H$,
@@
As we assumed a Fermi-like acceleration process, $\dot{E}=\alpha_{\rm acc}E$, with $\alpha_{\rm acc}$ independent of energy, thus
@@
so the left-hand side is the non-thermal energy density $U_{\rm nt}$ by definition.
@@
where we used $c_{s}=H\Omega$ and $U_{\rm th}=\tfrac{3}{2}P_{\rm th}$ with $\tfrac{5}{3}P_{\rm th}=\rho c_{s}^{2}$, as the disc is composed of mono-atomic gas in the acceleration region. Substituting into Eq.(15) gives,
@@
This condition fixes the normalisation. Hence the injected number density $n_{\rm inj}$ of the distribution of Eq.(8) is,
@@
where $L_{r}/H$ follows Eq.(2), $\beta=1/(\alpha_{\rm acc}\tau_{\rm esc})=1.5$ from Eq.(5), $M_{A}$ is computed locally (Appendix A), $\epsilon_{\rm inj}$ is obtained by numerically solving Eq. (12), $\gamma$ is likewise solved numerically and $\kappa$ remains the free energetics parameter.
Using the injected non-thermal density from Eq.(20), we compute the local non-thermal energy density $U_{\rm nt}(R,Z)$. Because the thermochemical background, thus the acceleration parameters, vary with the position, all quantities are evaluated cell by cell on the ProDiMo (R,Z) grid. Fig. 3 maps the resulting $U_{\rm nt}/U_{\rm th}$ and marks, in red, the regions where the local maximum particle energy exceeds $10\mathrm{MeV}$ (efficient acceleration  [^5] By efficient we mean, particles with energies above the peak of ionisation cross section.), while the black contour traces $N=10^{19}\mathrm{cm^{-2}}$ (the disc surface), we set the non-thermal power to $\kappa=0.1$, by analogy with the canonical $\sim 10-30\%$ efficiency with which supernova-remnant shocks are thought to channel kinetic energy into GCR (Ptuskin et al., 2010).
On Fig. 3, the sharp drop from $U_{\rm nt}>0$ to $U_{\rm nt}=0$ on the disc side tracks the $M_{A}=1$ contour. In our prescription, the maximum non-thermal energy $\varepsilon_{\max}$ tends to unity for $M_{A}\gtrsim 1$ (Eq.(9)), the supra-thermal tail merges with the thermal distribution. We therefore set $U_{\mathrm{nt}}=0$ for $M_{A}\gtrsim 1$, noting that this reflects the disappearance of a distinct non-thermal particle population rather than a suppression of turbulent reconnection itself. Moreover, because the turbulent reconnection model is formulated for strongly magnetised, low-$\beta_{p}$ plasmas (Lazarian and Vishniac, 1999), and simulations indicate that as $\beta_{p}$ increases, the non-thermal particle acceleration efficiency declines (Li et al., 2017). We restrict EP acceleration to low-$\beta_{p}$ regions and set $U_{\mathrm{nt}}=0$ for $\beta>10$. We note that efficient non-thermal acceleration may occur in hot or relativistic high-$\beta$ plasmas, yet such regimes are not relevant for protoplanetary discs and therefore lie outside the scope of this work.
@@
Notably, the map shows that our turbulence-driven reconnection model can accelerate particles to MeV energies in extended disc region, out to a few tens of AU along the ionised atmospheric layers, i.e. within the intermediate window where $M_{A}<1$ yet not so small that $\Delta$ allow $E_{\max}\gg E_{\rm th}$.
Throughout the acceleration region, $U_{\rm nt}$ is orders of magnitude smaller than the thermal energy density $U_{th}$ and, since $\beta_{p}\lesssim 1$, also far below the magnetic energy density. Consequently, EPs are not expected to exert a dynamical back-reaction over the plasma. This validates our treatment in the next section where EP transport is computed in the test-particle limit as a first approximation.
Figure 3: Non-thermal–to–thermal energy density ratio, $U_{\rm nt}/U_{\rm th}$ for $\kappa=0.1$. The red contour encloses regions where the local maximum particle energy exceeds $10\penalty 10000\ \mathrm{MeV}$ (efficient acceleration sites capable of supplying ionising EPs). The black contour marks a vertical column density $N=10^{19}\mathrm{cm^{-2}}$ (approximate disc surface).
@@
4 Results – Ionisation from reconnection-accelerated particles
To quantify the ionisation rate in the disc due to EPs generated by turbulence induced magnetic reconnection, we first compute their initial energy distribution, propagation and energy attenuation through the disc material.
@@ 4.1 Non-thermal particle energy distributions: injection and propagation
For each ProDiMo –R,Z grid cell, the local initial (”injection”) non-thermal energy distribution $F^{R,Z}_{\rm nt}(E)$ is computed with Eq. (8) at time $t=\tau_{\rm CS}$, using the normalisation of Eq. (20), together with the acceleration parameters $\alpha_{\rm acc}$, $\tau_{\rm esc}$ and $\tau_{\rm inj}$.
Fig. 4 shows representative energy distributions at $R=1\penalty 10000\ \mathrm{AU}$ within the height range where $U_{\rm nt}>0$: the thermal Maxwellian (dashed) based on ProDiMo density/temperature, and the associated non-thermal tail (solid) produced by our model of magnetic reconnection in turbulence.
@@
We then propagate this local energy distribution through the disc material with the continuous–slowing–down approximation (CSDA) to obtain, for any traversed column density $N$, the propagated distributions $F^{R,Z}_{\rm nt}(E,N)$ for both electrons and protons, as in Brunn et al. (2023). We assume isotropic injection and approximate trajectories as vertical, which is adequate to first order in the subsonic–to–transonic turbulent regime considered here. A more complete treatment that explicitly includes turbulence, accounting for spatial diffusion along turbulent field lines and for stochastic re-acceleration in MRI turbulence (e.g. Sun and Bai, 2021), is deferred to a forthcoming paper. Methodological details on loss functions, cross sections, CSDA validity, and the treatment of primaries/secondaries are provided by Padovani et al. (2009, 2018) and Brunn et al. (2023).
@@ 4.2 Ionisation rate
Energetic electrons and protons distribution, produced in the turbulence-induced magnetic reconnection, propagate into the protoplanetary disc and interact with the ambient gas, leading to the ionisation of hydrogen and molecular hydrogen. To compute the ionisation rate, we consider both primary ionisation reactions and secondary processes [^6] There is a typographical error in the expression of the differential ionisation cross section for electron in Eq. 19 of Kim et al. (2000). The right-hand side of the equation corresponds to $d\sigma/dw$, not $d\sigma/dW$. Confusing these two variables can introduce an error of a factor 13.6 in the differential cross section. triggered by the ejected energetic electrons. The ionisation rate, $\zeta_{R,Z}(N)$ at a vertical column density $N$ produced by the EPs distribution $F^{R,Z}_{\rm nt}(E)$ originating from a pixel at position $R,Z$, is computed as in Brunn et al. (2023).
The total EP-induced ionisation rate at radius $R$, $\zeta_{R}(N)$, is computed by vertically averaging the contributions from all reconnection-active pixels (cells with non-zero non-thermal energy density; the coloured pixels in Fig. 3). Each acceleration site at height $Z_{i}$ within the radial slice contributes by $\zeta_{R,Z_{i}}(N)$ to the total propagated ionisation profile. The total ionisation is,
@@
We adopt a thickness-weighted vertical average because, at a fixed radius, multiple acceleration sites at different heights contribute simultaneously, and not equally, to the ionisation at depth $N$. Weighting each pixel’s contribution $\zeta_{R,Z_{i}}(N)$ by its vertical extent $\Delta Z_{i}$, and normalising by the total active thickness $\Delta Z$, approximates the continuous integral $\Delta Z^{-1}\int\zeta_{R,Z}(N)dZ$. This ensures that layers occupying a larger fraction of the reconnecting zone contribute proportionally more, eliminates biases due to grid resolution and provides a surrogate for the time/ensemble average of inherently patchy, intermittent reconnection. We ignore EP ionisation from the opposite disc side because the propagated flux is suppressed before reaching the mid-plane at column density $N\gtrsim 10^{25}\,\mathrm{cm^{-2}}$.
Fig. 5 illustrates the result at $R=1\penalty 10000\ \mathrm{AU}$, for $\kappa=0.1$. Coloured curves (shown with the “viridis” colour scale) are the individual $\zeta_{R,Z_{i}}(N)$ for EPs injected at different heights, the thick orange curve is the vertically averaged total from Eq. (21), and the black curve shows, for comparison, the stellar X-ray ionisation of a typical T Tauri object with an X-ray luminosity $L_{X}=10^{30}$ erg s-1. While X-rays dominate at low column density ($10^{16}-10^{23}\mathrm{cm^{-2}}$), EPs substantially enhance and exceed the ionisation at higher columns (above a few $10^{23}\mathrm{cm^{-2}}$). At column density $N\gtrsim 10^{25}\,\mathrm{cm^{-2}}$ (not shown in Fig. 5), the EP ionisation rate drop, and the GCR component overtakes close to the disc mid-plain. The column density at which EPs, X-rays or GCR dominate ionisation depends on the distance to the star; we discuss this aspect in next section.
Figure 5: Ionisation rates as function of the disc vertical column density at $R=1$ AU for $\kappa=0.1$. The lines in “viridis” colour scale are the contribution to the ionisation rate coming from EPs accelerated in region at different $Z/R$. The thick orange line is the total ionisation rate corresponding to the weighted sum (Eq. 21) of each local contribution. The black line is the ionisation rate from the stellar X-rays.
@@ 4.3 Spatial ionisation rate distribution
Using Eq. (21), we compute the total EP–induced ionisation rate at every ProDiMo grid cell $(R,Z)$, yielding the map in Fig. 6. The figure shows $\zeta_{\rm EP}$ on a logarithmic colour scale from $10^{-20}$s-1 (dark purple) to $10^{-8}\mathrm{s^{-1}}$ (bright yellow). Dashed black curves are isocontours.
@@ 4.3.1 EP-dominated ionisation regions
Figure 7: Ionisation rate due to EPs ($\zeta_{\text{EP}}$) compared to the total ionisation rate $\zeta_{\rm tot}=\zeta_{\text{XR}}+\zeta_{\text{GCR}}+\zeta_{\text{EP}}$ for the reference case $\kappa=0.1$, where $\zeta_{\text{XR}}$ and $\zeta_{\text{GCR}}$ are the ionisation rates due to X-ray and Galactic Cosmic Ray, respectfully. The solid black contours mark regions where ionisation from energetic particles dominates over all other ionisation sources while the dashed black line indicates where the EP ionisation is 5% of the total ionisation rate. The red contour outlines regions dominated by Galactic Cosmic Rays. The region in between is dominated by X-ray ionisation.
@@
Fig. 7 shows the fraction of the total ionisation rate produced by EPs, $\zeta_{\rm EP}/\zeta_{\rm tot}$, for our reference case $\kappa=0.1$, with $\zeta_{\rm tot}=\zeta_{\rm EP}+\zeta_{\rm XR}+\zeta_{\rm GCR}$. The colour scale is logarithmic. Bright yellow marks regions where EPs contribute most. Solid black contours enclose zones where EPs dominate, i.e. $\zeta_{\rm EP}>\zeta_{\rm XR}+\zeta_{\rm GCR}$. The dashed black contour marks where EPs supply a non-negligible share of the ionisation rate, at least 5% of $\zeta_{\rm tot}$. The red contour outlines regions dominated by GCRs. Fig. 7 highlights the efficiency of EP ionisation even when only 10% of the local viscous dissipation is channelled into non-thermal particles. Even with this modest energy budget, EPs form a clear dominant band in the inner disc, between 0.120.1-2 AU in the intermediate/deep layer at $Z/R\sim 0.1$. Beyond the strictly dominant zones, large areas exceed the 5% threshold at distances between 0.1200.1-20 AU from the star, in the disc atmosphere and in the intermediate/deep layer between $Z/R\sim 0.8-0.15$. These zones of EP dominance are largely determined by the EP energy and propagation depth. In particular, GCRs due to their higher energies, propagate at higher column density and dominate the ionisation close the disc mid-plane. While X-rays dominate in the surface/intermediate region, above the EP-dominated zone. EPs are thus a significant contributor to the ionisation rate over extended regions, relevant for disc chemistry and dynamics.
Figure 8: Same as Fig. 7 for $\kappa=1$.
@@
Fig. 8 shows the spatial distribution of the EP contribution to the total ionisation rate under the extreme assumption that all local viscous dissipation is channelled to non-thermal particles, i.e. $\kappa=1$. The ionisation map shows that EPs dominate in the disc atmosphere from $\sim 0.2-40$ AU, and also in intermediate/deep layers from $\sim 0.08-20$ AU. This figure highlights that EPs accelerated by turbulence-induced magnetic reconnection can dominate the ionisation in several regions of the disc, and if not dominant, contribute significantly (more than 5%) to the total ionisation rate in most of the disc at radii from 0.1 AU to tens of AU.
@@
The $\kappa=1$ assumption is unrealistic, but we use it to illustrate the upper-limit spatial pattern that our model can produce. We note that a similar ionisation state could also occur in systems with higher accretion rates than adopted here. The value $\dot{M}=10^{-8}M_{\odot}$ yr-1 is a conservative choice. A larger $\dot{M}$ increases the viscous dissipation power and therefore the energy available for EP acceleration. In this sense, the $\kappa=1$ map can be viewed either as an upper bound or as a proxy for cases with $\kappa<1$ but higher $\dot{M}$, which can produce EP-dominated regions of comparable extent and depth.
The above results demonstrate that EPs, previously unexplored, may be a major ionisation source in thermochemical models of protoplanetary discs and could be relevant in planetary astrophysics and prebiotic chemistry.

### v2

@@ 3.3 Normalisation of the energy distribution from the available energy of the turbulent accretion
We normalised the non-thermal distribution by assuming that the power sustaining it, $\dot{U}_{\rm nt}$, is a fixed fraction, $\kappa$, of the local viscous dissipation per unit volume, $D/H$,
@@
As we assumed a Fermi-like acceleration process, $\dot{E}=\alpha_{\rm acc}E$, with $\alpha_{\rm acc}$ independent of energy, and thus
@@
so the left-hand side is the non-thermal energy density, $U_{\rm nt}$, by definition.
@@
where we used $c_{s}=H\Omega$ and $U_{\rm th}=\tfrac{3}{2}P_{\rm th}$ with $\tfrac{5}{3}P_{\rm th}=\rho c_{s}^{2}$, as the disc is composed of mono-atomic gas in the acceleration region. Substituting into Eq.(15) gives
@@
This condition fixes the normalisation. Hence the injected number density, $n_{\rm inj}$, of the distribution of Eq.(8) is
@@
where $L_{r}/H$ follows Eq.(2), $\beta=1/(\alpha_{\rm acc}\tau_{\rm esc})=1.5$ from Eq.(5), $M_{A}$ is computed locally (Appendix A), $\epsilon_{\rm inj}$ is obtained by numerically solving Eq. (12), $\gamma$ is likewise solved numerically, and $\kappa$ remains the free energetics parameter.
Using the injected non-thermal density from Eq.(20), we computed the local non-thermal energy density, $U_{\rm nt}(R,Z)$. Because the thermochemical background, and thus the acceleration parameters, vary with the position, all quantities were evaluated cell by cell on the ProDiMo (R,Z) grid. Fig. 3 maps the resulting $U_{\rm nt}/U_{\rm th}$ and marks, in red, the regions where the local maximum particle energy exceeds $10\mathrm{MeV}$ (efficient acceleration  [^5] By efficient we mean, particles with energies above the peak of ionisation cross section.), while the black contour traces $N=10^{19}\mathrm{cm^{-2}}$ (the disc surface). We set the non-thermal power to $\kappa=0.1$, by analogy with the canonical $\sim 10-30\%$ efficiency with which supernova-remnant shocks are thought to channel kinetic energy into GCR (Ptuskin et al., 2010).
On Fig. 3, the sharp drop from $U_{\rm nt}>0$ to $U_{\rm nt}=0$ on the disc side tracks the $M_{A}=1$ contour. In our prescription, the maximum non-thermal energy $\varepsilon_{\max}$ tends to unity for $M_{A}\gtrsim 1$ (Eq.(9)), the supra-thermal tail merges with the thermal distribution. We therefore set $U_{\mathrm{nt}}=0$ for $M_{A}\gtrsim 1$, noting that this reflects the disappearance of a distinct non-thermal particle population rather than a suppression of turbulent reconnection itself. Moreover, because the turbulent reconnection model is formulated for strongly magnetised, low-$\beta_{p}$ plasmas (Lazarian and Vishniac, 1999), and simulations indicate that as $\beta_{p}$ increases, the non-thermal particle acceleration efficiency declines (Li et al., 2017). We restricted EP acceleration to low-$\beta_{p}$ regions and set $U_{\mathrm{nt}}=0$ for $\beta>10$. We note that efficient non-thermal acceleration may occur in hot or relativistic high-$\beta$ plasmas, yet such regimes are not relevant for protoplanetary discs, and therefore lie outside the scope of this work.
@@
Notably, the map shows that our turbulence-driven reconnection model can accelerate particles to mega-electronvolt energies in extended disc region, out to a few tens of astronomical units along the ionised atmospheric layers, i.e. within the intermediate window where $M_{A}<1$ yet not so small that $\Delta$ allow $E_{\max}\gg E_{\rm th}$.
Throughout the acceleration region, $U_{\rm nt}$ is orders of magnitude smaller than the thermal energy density, $U_{th}$, and, since $\beta_{p}\lesssim 1$, also far below the magnetic energy density. Consequently, EPs are not expected to exert a dynamical back-reaction over the plasma. This validates our treatment in the next section where EP transport is computed in the test-particle limit as a first approximation.
Figure 3: Non-thermal–to–thermal energy density ratio, $U_{\rm nt}/U_{\rm th}$, for $\kappa=0.1$. The red contour encloses regions where the local maximum particle energy exceeds $10\penalty 10000\ \mathrm{MeV}$ (efficient acceleration sites capable of supplying ionising EPs). The black contour marks a vertical column density, $N=10^{19}\mathrm{cm^{-2}}$ (approximate disc surface).
@@
4 Results: Ionisation from reconnection-accelerated particles
To quantify the ionisation rate in the disc due to EPs generated by turbulence induced magnetic reconnection, we first computed their initial energy distribution, propagation and energy attenuation through the disc material.
@@ 4.1 Non-thermal particle energy distributions: injection and propagation
For each ProDiMo –R,Z grid cell, the local initial (’injection’) non-thermal energy distribution $F^{R,Z}_{\rm nt}(E)$, was computed with Eq. (8) at time $t=\tau_{\rm CS}$, using the normalisation of Eq. (20), together with the acceleration parameters $\alpha_{\rm acc}$, $\tau_{\rm esc}$, and $\tau_{\rm inj}$. Fig. 4 shows representative energy distributions at $R=1\penalty 10000\ \mathrm{AU}$ within the height range where $U_{\rm nt}>0$: the thermal Maxwellian (dashed) based on ProDiMo density and temperature, and the associated non-thermal tail (solid) produced by our model of magnetic reconnection in turbulence.
@@
We then propagated this local energy distribution through the disc material with the continuous–slowing–down approximation (CSDA) to obtain, for any traversed column density, $N$, the propagated distributions, $F^{R,Z}_{\rm nt}(E,N)$, for both electrons and protons, as in Brunn et al. (2023). We assume isotropic injection and approximate trajectories as vertical, which is adequate to first order in the subsonic–to–transonic turbulent regime considered here. A more complete treatment that explicitly includes turbulence, accounting for spatial diffusion along turbulent field lines and for stochastic re-acceleration in MRI turbulence (e.g. Sun and Bai, 2021), is deferred to a forthcoming paper. Methodological details on loss functions, cross sections, CSDA validity, and the treatment of primaries and secondaries are provided by Padovani et al. (2009, 2018) and Brunn et al. (2023).
@@ 4.2 Ionisation rate
Energetic electrons and protons, produced in the turbulence-induced magnetic reconnection, propagate into the protoplanetary disc and interact with the ambient gas, leading to the ionisation of hydrogen and molecular hydrogen. To compute the ionisation rate, we consider both primary ionisation reactions and secondary processes [^6] There is a typographical error in the expression of the differential ionisation cross section for electron in Eq. 19 of Kim et al. (2000). The right-hand side of the equation corresponds to $d\sigma/dw$, not $d\sigma/dW$. Confusing these two variables can introduce an error of a factor 13.6 in the differential cross section. triggered by the ejected energetic electrons. The ionisation rate, $\zeta_{R,Z}(N)$, at a vertical column density, $N$, produced by the EP distribution, $F^{R,Z}_{\rm nt}(E)$, originating from a pixel at position $R,Z$, was computed as in Brunn et al. (2023).
The total EP-induced ionisation rate at radius $R$, $\zeta_{R}(N)$, was computed by vertically averaging the contributions from all reconnection-active pixels (cells with non-zero non-thermal energy density; the coloured pixels in Fig. 3). Each acceleration site at height $Z_{i}$ within the radial slice contributes by $\zeta_{R,Z_{i}}(N)$ to the total propagated ionisation profile. The total ionisation is
@@
We adopt a thickness-weighted vertical average because, at a fixed radius, multiple acceleration sites at different heights contribute simultaneously, and not equally, to the ionisation at depth $N$. Weighting each pixel’s contribution, $\zeta_{R,Z_{i}}(N)$, by its vertical extent, $\Delta Z_{i}$, and normalising by the total active thickness, $\Delta Z$, approximates the continuous integral, $\Delta Z^{-1}\int\zeta_{R,Z}(N)dZ$. This ensures that layers occupying a larger fraction of the reconnecting zone contribute proportionally more, eliminates biases due to grid resolution, and provides a surrogate for the time and ensemble average of inherently patchy, intermittent reconnection. We ignored EP ionisation from the opposite disc side because the propagated flux is suppressed before reaching the mid-plane at column density $N\gtrsim 10^{25}\,\mathrm{cm^{-2}}$.
Fig. 5 illustrates the result at $R=1\penalty 10000\ \mathrm{AU}$, for $\kappa=0.1$. Coloured curves (shown with the ’viridis’ colour scale) are the individual $\zeta_{R,Z_{i}}(N)$ for EPs injected at different heights, the thick orange curve is the vertically averaged total from Eq. (21), and the black curve shows, is for comparison, the stellar X-ray ionisation of a typical T Tauri object with an X-ray luminosity of $L_{X}=10^{30}$ erg s-1. While X-rays dominate at low column density ($10^{16}-10^{23}\mathrm{cm^{-2}}$), EPs substantially enhance and exceed the ionisation at higher columns (above a few $10^{23}\mathrm{cm^{-2}}$). At column density $N\gtrsim 10^{25}\,\mathrm{cm^{-2}}$ (not shown in Fig. 5), the EP ionisation rate drops, and the GCRs component overtakes close to the disc mid-plain. The column density at which EPs, X-rays or GCR dominate ionisation depends on the distance to the star; we discuss this aspect in the next section.
Figure 5: Ionisation rates as function of the disc vertical column density at $R=1$ AU for $\kappa=0.1$. The lines in the viridis colour scale are the contribution to the ionisation rate coming from EPs accelerated in region at different $Z/R$. The thick orange line is the total ionisation rate corresponding to the weighted sum (Eq. 21) of each local contribution. The black line is the ionisation rate from the stellar X-rays.
@@ 4.3 Spatial ionisation rate distribution
Using Eq. (21), we computed the total EP–induced ionisation rate at every ProDiMo grid cell $(R,Z)$, yielding the map in Fig. 6. The figure shows $\zeta_{\rm EP}$ on a logarithmic colour scale from $10^{-20}$s-1 (dark purple) to $10^{-8}\mathrm{s^{-1}}$ (bright yellow). Dashed black curves are isocontours.
@@ 4.3.1 EP-dominated ionisation regions
Figure 7: Ionisation rate due to EPs ($\zeta_{\text{EP}}$) compared to the total ionisation rate $\zeta_{\rm tot}=\zeta_{\text{XR}}+\zeta_{\text{GCR}}+\zeta_{\text{EP}}$, for the reference case $\kappa=0.1$, where $\zeta_{\text{XR}}$ and $\zeta_{\text{GCR}}$ are the ionisation rates due to X-rays and GCRs, respectfully. The solid black contours mark regions where ionisation from EPs dominates over all other ionisation sources, while the dashed black line indicates where the EP ionisation is 5% of the total ionisation rate. The red contour outlines regions dominated by GCRs. The region in between is dominated by X-ray ionisation.
@@
Fig. 7 shows the fraction of the total ionisation rate produced by EPs, $\zeta_{\rm EP}/\zeta_{\rm tot}$, for our reference case $\kappa=0.1$, with $\zeta_{\rm tot}=\zeta_{\rm EP}+\zeta_{\rm XR}+\zeta_{\rm GCR}$. The colour scale is logarithmic. Bright yellow marks regions where EPs contribute most. Solid black contours enclose zones where EPs dominate, i.e. $\zeta_{\rm EP}>\zeta_{\rm XR}+\zeta_{\rm GCR}$. The dashed black contour marks where EPs supply a non-negligible share of the ionisation rate, at least 5% of $\zeta_{\rm tot}$. The red contour outlines regions dominated by GCRs. Fig. 7 highlights the efficiency of EP ionisation even when only 10% of the local viscous dissipation is channelled into non-thermal particles. Even with this modest energy budget, EPs form a clear dominant band in the inner disc, between 0.120.1-2 AU in the intermediate to deep layer at $Z/R\sim 0.1$. Beyond the strictly dominant zones, large areas exceed the 5% threshold at distances between 0.1200.1-20 AU from the star, in the disc atmosphere and in the intermediate to deep layer between $Z/R\sim 0.8-0.15$. These zones of EP dominance are largely determined by the EP energy and propagation depth. In particular, GCRs due to their higher energies, propagate at higher column density and dominate the ionisation close the disc mid-plane. While X-rays dominate in the surface to intermediate region, above the EP-dominated zone. The EPs are thus a significant contributor to the ionisation rate over extended regions, relevant for disc chemistry and dynamics.
Figure 8: Same as Fig. 7, but for $\kappa=1$.
@@
Fig. 8 shows the spatial distribution of the EP contribution to the total ionisation rate under the extreme assumption that all local viscous dissipation is channelled to non-thermal particles, i.e. $\kappa=1$. The ionisation map shows that EPs dominate in the disc atmosphere from $\sim 0.2-40$ AU, and also in intermediate to deep layers from $\sim 0.08-20$ AU. This figure highlights that EPs accelerated by turbulence-induced magnetic reconnection can dominate the ionisation in several regions of the disc, and if not dominant, contribute significantly (more than 5%) to the total ionisation rate in most of the disc at radii from 0.1 AU to tens of astronomical units.
@@
The $\kappa=1$ assumption is unrealistic, but we use it to illustrate the upper-limit spatial pattern that our model can produce. We note that a similar ionisation state could also occur in systems with higher accretion rates than adopted here. The value $\dot{M}=10^{-8}M_{\odot}$ yr-1 is a conservative choice. A larger $\dot{M}$ increases the viscous dissipation power, and therefore the energy available for EP acceleration. In this sense, the $\kappa=1$ map can be viewed either as an upper bound or as a proxy for cases with $\kappa<1$ but higher $\dot{M}$, which can produce EP-dominated regions of comparable extent and depth. The above results demonstrate that EPs, previously unexplored, may be a major ionisation source in thermochemical models of protoplanetary discs and could be relevant in planetary astrophysics and prebiotic chemistry.